Simplify: (n+1)!/(n-1)!

I am really stuck on this problem and I just really need help with this.

Answer by ewatrrr(24785)


Hi
Simplify:
(n+1!/(n-1)! = (n+1)*n*(n-1)(n-2).../(n-1)*(n-2)... = (n+1)n OR n(n+1)
Whenever doing division problems with factorials,always look for the
results of cancelation of terms in the numerator by those in the denominator.